Write down expectation value of r^2for hydrogen atom.
1
Expert's answer
2020-11-09T10:14:59-0500

Expectation value of "r^2":


"\big<r^2\big>=\int^\infty_0\psi_{1s}(r)r^2\psi_{1s}(r)4\pi r^2\text{d}r=3."